# CrystalBox: Future-Based Explanations for Input-Driven Deep RL Systems

### v3

@@ 1 Introduction
(a) (a) CrystalBox’s explanation for $S_{1}$ showing that medium quality action provides a higher future return due to it lowering stalls.
(b) (b) CrystalBox’s explanation for $S_{2}$ showing that high-quality action provides a higher return due to the quality component.
Figure 2: CrystalBox’s explanation for the two motivating states in Section 2. CrystalBox allows us to quickly understand why the controller’s actions are more appropriate in both states by letting us compare their decomposed future returns to those of alternative actions.
@@ 2 Motivation
Feature-based approach. We choose LIME, a widely employed explainer (Ribeiro, Singh, and Guestrin 2016), as a representative feature-based explainer and discuss its ability to answer $Q1$ and $Q2$. LIME is a model-agnostic, local explainer that creates a linear explanation pinpointing the top features driving a given action from an input state. Figure 0(b) shows the explanations generated by LIME in the state $S_{1}$ for the two actions medium and high quality bitrates, and 0(c) shows the same for the state $S_{2}$.
@@ 4 Future-Based Explanations
Figure 3: System Diagram of CrystalBox: CrystalBox consists of two components: a learned decomposed returns predictor and a post-processing module. We train a function approximator once to predict the decomposed returns by (i) collecting rollouts of the policy in the simulation environment, pre-processing the rollouts to form a dataset, and (ii) employing supervised learning. Once trained, we give the query state and action to this approximator, obtain its predicted decomposed returns, and optionally post-process them to generate explanations.
@@ 5 CrystalBox
First, we propose a procedure to obtain samples of $Q^{\pi}_{c}$. From Eq. 2, we see that $Q^{\pi}_{c}$ depends on the combined transition function of the environment. This combined transition function forms a power distribution involving the variables $P_{z}$ and $P_{s}$. $P_{z}$ cannot be approximated with low bias or variance, leading to high bias estimations of $Q^{\pi}_{c}$. To solve this problem, we circumvent approximating $P_{z}$ by instead using its unbiased samples. These unbiased samples are exactly the traces from the training set of traces $Z$. We replay these traces with a given policy $\pi$ to obtain estimates of $Q^{\pi}_{c}$ (See Fig. 2(b) ‘Data Collection’ block) using Monte-Carlo rollouts (Sutton and Barto 2018). Second, we define our function approximator architecture with parameters $\theta$. We use a simple split fully connected layers architecture (Appendix A.5111Appendices can be found in the extended version of this paper (Patel, Jyothi, and Narodytska 2023)) to jointly learn the return components. Finally, we define a loss for each individual component $c$ and take a weighted sum as the overall loss. We use standard Mean Squared Error for our individual components.
@@ 6.2 Sampling Techniques
Naive Sampling A simple strategy for sampling involves uniformly random sampling $z$ from our training set $Z$. We observed that this method suffers from low accuracy (see § 7) due to its dependence on the dataset distribution, which can be unbalanced in input-driven applications. Dominant traces often do not cover all relevant scenarios (see Fig. 11 and 12 in Appendix A.1).
@@ 7.2 Fidelity Evaluation
In Figure 4, we show the error of returns predicted by CrystalBox and sampling baselines. We see that CrystalBox outperforms both of the sampling approaches in generating high-fidelity predictions for all three return components across both environments, under factual (Fig. 3(a)3(c)) and counterfactual (Fig. 3(b)3(d)) actions. Note that despite ABR’s discrete action space and CC’s continuous one, CrystalBox offers high-fidelity explanations in each case. We believe that this is due to CrystalBox’s estimates being unbiased by construction, unlike the ones from sampling baselines.
@@ 7.3 CrystalBox Analysis
(a) Figure 4: Tuning the weight of Stalling Reward Component in ABR. Here, we employ CrystalBox to explain why the controller chooses to drop its bitrate in seemingly good states. We identify the dominant reward component and plot the distribution.
@@
(b) (a) ABR: Recall
(c) (b) ABR: False Positive Rate
(d) (c) CC: Recall
(e) (d) CC: False Positive Rate
@@ 7.4 CrystalBox Applications
Figure 6 shows our results for ABR and CC under factual actions. Observing the ABR results, we note the following. CrystalBox exhibits high recall and low false-positive rates for both quality drops and long stalls in factual explanations. Sampling-based methods miss all long stalls but detect large quality changes better, albeit with higher false-positives. These observations hold for the CC environment. Similar findings for counterfactual actions are in Appendix A.2.
@@ 8 Discussion and Future Work
Investigating the use of feature-based techniques for creating an interpretable future return predictors represents a compelling research direction. Another potential avenue is to explore whether we can employ future return predictors during policy learning for human-in-the-loop frameworks.
@@ A.5 CrystalBox Details
For the fully connected layers, we perform limited tuning to choose the units of these layers from {64, 128, 256, 512}. We found that a smaller number of units is enough in both of our environments. We present a visualization of our architectures in Figures 13 and 14.

### v4

@@ 1 Introduction
(a) (a) CrystalBox’s explanation for $S_{1}$.
(b) (b) CrystalBox’s explanation for $S_{2}$.
Figure 2: CrystalBox’s explanation for the two motivating states in Section 2. CrystalBox allows us to understand why the controller’s actions are appropriate in both states by letting us compare their decomposed future returns to those of alternatives.
@@ 2 Motivation
Feature-based approach. We choose LIME, a widely employed explainer (Ribeiro, Singh, and Guestrin 2016), as a representative feature-based explainer and discuss its ability to answer $Q1$ and $Q2$. LIME is a model-agnostic, local explainer that creates a linear explanation pinpointing the top features driving a given action from an input state. Figure 1(b) shows the explanations generated by LIME in the state $S_{1}$ for the two actions medium and high quality bitrates, and 1(c) shows the same for the state $S_{2}$.
@@ 4 Future-Based Explanations
Figure 3: System Diagram of CrystalBox: CrystalBox consists of two components: a learned decomposed returns predictor and a post-processing module. We train a function approximator once to predict the decomposed returns by (i) collecting rollouts of the policy, pre-processing the rollouts to form a dataset, and (ii) employing supervised learning. Once trained, we give the query state and action to this approximator, obtain its predicted decomposed returns, and optionally post-process them.
@@ 5 CrystalBox
First, we propose a procedure to obtain samples of $Q^{\pi}_{c}$. From Eq. 2, we see that $Q^{\pi}_{c}$ depends on the combined transition function of the environment. This combined transition function forms a power distribution involving the variables $P_{z}$ and $P_{s}$. $P_{z}$ cannot be approximated with low bias or variance, leading to high bias estimations of $Q^{\pi}_{c}$. To solve this problem, we circumvent approximating $P_{z}$ by instead using its unbiased samples. These unbiased samples are exactly the traces from the training set of traces $Z$. We replay these traces with a given policy $\pi$ to obtain estimates of $Q^{\pi}_{c}$ (See Fig. 3(b) ‘Data Collection’ block) using Monte-Carlo rollouts (Sutton and Barto 2018). Second, we define our function approximator architecture with parameters $\theta$. We use a simple split fully connected layers architecture (Appendix A.5111This is the extended version of the paper (Patel, Abdu Jyothi, and Narodytska 2024)) to jointly learn the return components. Finally, we define a loss for each individual component $c$ and take a weighted sum as the overall loss. We use standard Mean Squared Error for our individual components.
@@ 6.2 Sampling Techniques
Naive Sampling A simple strategy for sampling involves uniformly random sampling $z$ from our training set $Z$. We observed that this method suffers from low accuracy (see § 7) due to its dependence on the dataset distribution, which can be unbalanced in input-driven applications. Dominant traces often do not cover all relevant scenarios (see Fig. 12 and 13 in Appendix A.1).
@@ 7.2 Fidelity Evaluation
In Figure 4, we show the error of returns predicted by CrystalBox and sampling baselines. We see that CrystalBox outperforms both of the sampling approaches in generating high-fidelity predictions for all three return components across both environments, under factual (Fig. 4(a)4(c)) and counterfactual (Fig. 4(b)4(d)) actions. Note that despite ABR’s discrete action space and CC’s continuous one, CrystalBox offers high-fidelity explanations in each case. We believe that this is due to CrystalBox’s estimates being unbiased by construction, unlike the ones from sampling baselines.
@@ 7.3 CrystalBox Analysis
(a) Figure 5: Tuning the weight of Stalling Reward Component in ABR. Here, we employ CrystalBox to explain why the controller chooses to drop bitrate in seemingly good states. We identify the dominant reward components and plot them.
@@
(a) (a) ABR: Recall
(b) (b) ABR: False Positive Rate
(c) (c) CC: Recall
(d) (d) CC: False Positive Rate
@@ 7.4 CrystalBox Applications
Figure 7 shows our results for ABR and CC under factual actions. Observing the ABR results, we note the following. CrystalBox exhibits high recall and low false-positive rates for both quality drops and long stalls in factual explanations. Sampling-based methods miss all long stalls but detect large quality changes better, albeit with higher false positives. These observations hold for the CC environment. Similar findings for counterfactual actions are in Appendix A.2.
@@ 8 Discussion and Future Work
Investigating the use of feature-based techniques for creating interpretable future return predictors represents a compelling research direction. Another potential avenue is to explore whether we can employ future return predictors during policy learning for human-in-the-loop frameworks.
@@ A.5 CrystalBox Details
For the fully connected layers, we perform limited tuning to choose the units of these layers from {64, 128, 256, 512}. We found that a smaller number of units is enough in both of our environments. We present a visualization of our architectures in Figures 14 and 15.